# GS4: Generalizable Sparse Splatting Semantic SLAM

### v2

@@ 4.2 Results
Rendering and Reconstruction Performance. In Table 1, we evaluate the rendering and reconstruction performance of our method on ScanNet. This is a difficult task compared to the synthetic data where neural RGB-D SLAM methods usually show strong results, because inevitably inaccurate ground truth camera poses and depths make optimization much harder than completely clean synthetic datasets. Compared to existing dense neural RGB-D SLAM methods, our approach achieves state-of-the-art performance on PSNR, SSIM, and Depth L1 metrics. Specifically, our method surpasses the runner-up, GS3LAM (Li et al., 2023), by 3.59 dB in PSNR (a 17.4 percent improvement). and 0.091 in SSIM (a 10.4 percent improvement). Furthermore, our approach utilizes approximately 10x fewer Gaussians than the baselines. This efficiency highlights the effectiveness of our method in achieving high-quality scene representation with reduced computational complexity.
@@
| GS4 (Ours, AFF, 1 iter) | 22.61 | 0.850 | 0.335 | 6.55 | 356k | |
| GS4 (Ours, AFF, 5 iters) | 24.26 | 0.879 | 0.304 | 4.98 | 245k | |
@@
| GS4 (Ours, AFF, 1 iter) | 22.50 | 0.885 | 0.238 | 6.04 | 172k | |
| GS4 (Ours, AFF, 5 iters) | 24.02 | 0.915 | 0.201 | 5.05 | 130k |
@@
Semantic Performance. In Table 2, we present both 2D rendering and 3D mean Intersection over Union (mIoU) scores across the six ScanNet test scenes. For 3D mIoU evaluation, we first align the reconstructed map with the ground-truth mesh and then use 3D neighborhood voting to assign predicted labels. Our method outperforms the previous runner-up, GS3LAM, by 20.42 in 3D mIoU and by 6.62 in 2D mIoU. Qualitative comparisons are provided in the appendix.
| Methods | SGS SLAM | OVO-Gaussian-SLAM | GS3LAM | GS4 (Ours, AFF, 1 iter) | GS4 (Ours, AFF, 5 iters) |
| mIoU(2D) | 37.20 | ✗ | 56.42 | 63.04 | 62.36 |
@@
| Metric | NICE-SLAM | Point-SLAM | SplaTAM | RTG SLAM | GS-ICP SLAM | SGS SLAM | GS3LAM | GO-SLAM | GS4 (Ours) |
| ATE RMSE [cm] | 10.70 | 12.19 | 11.88 | 144.52 | NaN | 40.97 | 30.88 | 7.00 | 6.98 |
Runtime Comparison. Table 4 presents a runtime comparison of our method against the baselines at the resolution, conducted on an Nvidia RTX TITAN. FPS is calculated by dividing the total number of frames by the total time to represent the overall system performance. While GS-ICP SLAM is faster than ours, its rendering and tracking performance is significantly worse (Table 1 and 3). Our approach is 12x faster than SplaTAM, 17x faster than SGS-SLAM, and 24x faster than GS3LAM.
| Methods | Point-SLAM | SplaTAM | RTG-SLAM | GS-ICP SLAM | SGS-SLAM | GS3LAM | GS4 (ours, AFF, 1 iter) | GS4 (ours, AFF, 5 iters) |
| FPS | 0.05 | 0.23 | 1.01 | 3.62 | 0.17 | 0.12 | 2.87 | 1.92 |
ScanNet++ Experiments. Table 5 presents the 2D rendering results for both training and novel views, as well as the 3D semantic, tracking, and runtime comparisons of our method against the baselines at a resolution of 256384. The results are averaged over six scenes. Our method achieves state-of-the-art performance in both 2D and 3D tasks, showing a 10 improvement in PSNR and SSIM on the training view and a 13.22 improvement in 3D mIoU over the runner-up, while using only 2.0–2.5 of the number of Gaussians required by other methods. Compared to SGS-SLAM, although it achieves competitive 2D mIoU on the training view, its performance drops drastically on the novel view (2.46) and in 3D mIoU (0.58), indicating severe overfitting to the input views and poor generalization to the full 3D scene. In addition to delivering superior performance across both 2D and 3D metrics, our method also exhibits substantially higher efficiency, running 19× faster than SplaTAM and 25× faster than SGS-SLAM.
| Method | Train View | Novel View | mIoU (3D) | GS Num | ATE RMSE | FPS | ||||||||
| PSNR | SSIM | LPIPS | Depth L1 | mIoU (2D) | PSNR | SSIM | LPIPS | Depth L1 | mIoU (2D) | |||||
| SplaTAM | 22.98 | 0.854 | 0.218 | 7.11 | ✗ | 17.70 | 0.662 | 0.444 | 10.28 | ✗ | ✗ | 1899k | 1588.95 | 0.17 |
| SGS-SLAM | 21.29 | 0.801 | 0.276 | 10.52 | 28.24 | 9.70 | 0.265 | 0.664 | 86.93 | 2.46 | 0.58 | 2374k | 1315.89 | 0.13 |
| GS4 (Ours) | 25.26 | 0.929 | 0.162 | 21.93 | 23.51 | 15.61 | 0.531 | 0.425 | 31.90 | 13.27 | 13.80 | 48k | 5.88 | 3.23 |
Zero-shot Experiments. In Table 6, we report quantitative zero-shot results. For NYUv2, the numbers are averaged over three scenes, and for TUM-RGBD, they are also averaged over three scenes. Per-scene results are provided in the appendix. On NYUv2, our method outperforms all other GS-based SLAM approaches across all rendering metrics while using significantly fewer Gaussians. Qualitative comparisons are also provided in the appendix. On TUM-RGBD, our method outperforms the baselines in terms of SSIM and the number of Gaussians, and closely matches the best performance in other metrics, despite relying primarily on a feed-forward model trained on ScanNet.
@@
| GS4 (Ours, AFF, 5 iters) | 22.70 | 0.903 | 0.191 | 166k |
Ablation Study. We conduct an ablation study using all ScanNet test scenes, as shown in Table 7. The results demonstrate that both the Gaussian Refinement Network and the Few-Iteration Gaussian Optimization are critical to the performance of GS4. Additionally, Gaussian pruning significantly reduces the number of Gaussians without sacrificing accuracy. More results on backbone comparisons, reconstruction accuracy and qualitative results are shown in the appendix.
| Design Choice | PSNR [dB] | SSIM | LPIPS | Depth L1 | mIoU | Gs Num |
| GS Prediction | 14.91 | 0.460 | 0.663 | 33.14 | 41.0 | 133k |
| + GS Refinement | 16.15 | 0.556 | 0.584 | 29.56 | 44.68 | 576k |
| + Few-Iter. Optimization (1) | 22.66 | 0.852 | 0.335 | 6.41 | 63.7 | 680k |
| + GS Pruning (Full SLAM) | 22.61 | 0.851 | 0.335 | 6.56 | 63.1 | 355k |
@@ 5 Conclusion
We present GS4, a novel SLAM system that incrementally constructs and updates a 3D semantic scene representation from a monocular RGB-D video with a learned generalizable network. Our novel Gaussian refinement network and few-iteration Gaussian optimization significantly improve the performance of our approach. Our experiments demonstrate state-of-the-art semantic SLAM performance on the ScanNet benchmark while running 10x faster and using 10x less Gaussians than baselines. The model also showed strong generalization capabilities through zero-shot transfer to the NYUv2 and TUM RGB-D datasets. In future work, we will further improve the computational speed of GS4 and explore options for a pure RGB-based SLAM approach.
@@ Appendix A More Experimental Setup
For ScanNet, we evaluate on six scenes (scene0000, scene0059, scene0106, scene0169, scene0181, and scene0207), which are commonly used by other SLAM methods. For ScanNet++, we use six scenes in total: 8b5caf3398, 3f15a9266d, e7af285f7d, 99fa5c25e1, 09c1414f1b, and 9071e139d9.
Appendix B Reconstruction Results on ScanNet

### v3

@@ 4.2 Results
| GS4 (Ours, 1 iter) | 22.62 | 0.851 | 0.338 | 6.46 | 321k | |
| GS4 (Ours, 5 iters) | 24.55 | 0.885 | 0.299 | 4.86 | 224k | |
@@
| GS4 (Ours, 1 iter) | 22.54 | 0.885 | 0.240 | 5.98 | 162k | |
| GS4 (Ours, 5 iters) | 24.31 | 0.921 | 0.196 | 4.93 | 124k |
Rendering and Reconstruction Performance. In Table 1, we evaluate the rendering and reconstruction performance of our method on ScanNet. This is a difficult task compared to the synthetic data where neural RGB-D SLAM methods usually show strong results, because inevitably inaccurate ground truth camera poses and depths make optimization much harder than completely clean synthetic datasets. Compared to existing dense neural RGB-D SLAM methods, our approach achieves state-of-the-art performance on PSNR, SSIM, and Depth L1 metrics. Specifically, our method surpasses the runner-up, GS3LAM [li2023dnsslamdenseneural], by 3.88 dB in PSNR (a 18.8 percent improvement). and 0.089 in SSIM (a 11.2 percent improvement). Furthermore, our approach utilizes approximately 10x fewer Gaussians than the baselines. This efficiency highlights the effectiveness of our method in achieving high-quality scene representation with reduced computational complexity.
@@
Semantic Performance. In Table 2, we present both 2D rendering and 3D mean Intersection over Union (mIoU) scores across the six ScanNet test scenes. For 3D mIoU evaluation, we first align the reconstructed map with the ground-truth mesh and then use 3D neighborhood voting to assign predicted labels. Our method outperforms the previous runner-up, GS3LAM, by 20.42 in 3D mIoU and by 7.29 in 2D mIoU. Qualitative comparisons are provided in the appendix.
| Methods | SGS SLAM | OVO-Gaussian-SLAM | GS3LAM | GS4 (Ours, 1 iter) | GS4 (Ours, 5 iters) |
| mIoU(2D) | 37.20 | ✗ | 56.42 | 63.71 | 62.10 |
@@
| Metric | NICE-SLAM | Point-SLAM | SplaTAM | RTG SLAM | GS-ICP SLAM |
| ATE RMSE [cm] | 10.70 | 12.19 | 11.88 | 144.52 | NaN |
| Metric | SGS SLAM | GS3LAM | GO-SLAM | GS4 (Ours, 1 iter) | GS4 (Ours, 5 iters) |
| ATE RMSE [cm] | 40.97 | 30.88 | 7.00 | 6.97 | 6.95 |
| Methods | Point-SLAM | SplaTAM | RTG-SLAM | GS-ICP SLAM |
| FPS | 0.05 | 0.23 | 1.01 | 3.62 |
| Methods | SGS-SLAM | GS3LAM | GS4 (ours, 1 iter) | GS4 (ours, 5 iters) |
| FPS | 0.17 | 0.12 | 2.85 | 1.82 |
Runtime Comparison. Table 4 presents a runtime comparison of our method against the baselines at the resolution, conducted on an Nvidia RTX TITAN. FPS is calculated by dividing the total number of frames by the total time to represent the overall system performance. While GS-ICP SLAM is faster than ours, its rendering and tracking performance is significantly worse (Table 1 and 3). Our approach is 12x faster than SplaTAM, 17x faster than SGS-SLAM, and 24x faster than GS3LAM. Notably, even with this exceptional speed, our approach maintains superior map quality and outperforms other methods.
| Method | Rendering Metrics | mIoU (3D) | GS Num | ATE RMSE | FPS | |||
| PSNR | SSIM | LPIPS | mIoU (2D) | |||||
| SplaTAM | 20.34 | 0.758 | 0.331 | ✗ | ✗ | 1899k | 1588.95 | 0.17 |
| SGS-SLAM | 15.59 | 0.533 | 0.470 | 15.35 | 0.58 | 2374k | 1315.89 | 0.13 |
| GS4 (Ours) | 21.06 | 0.740 | 0.281 | 18.92 | 13.80 | 85k | 5.37 | 1.60 |
ScanNet++ Experiments. Table 5 reports 2D rendering performance averaged over both training and novel views, as well as 3D semantic, tracking, and runtime comparisons against the baselines at a resolution of 256384, averaged over six scenes. Our method achieves state-of-the-art performance on both 2D and 3D tasks, outperforming the runner-up by 0.72dB in PSNR (3.5 relative improvement), 0.05 in LPIPS (15.1 relative improvement), 3.57 in 2D mIoU, and 13.22 in 3D mIoU, while using only 3.0–4.5 of the Gaussians required by other methods. Compared to SGS-SLAM, although it achieves the second-best 2D mIoU, its 3D mIoU drops to just 0.58, indicating poor generalization to the full 3D scene. In addition to delivering superior performance across both 2D and 3D metrics, our method also exhibits substantially higher efficiency, running 9× faster than SplaTAM and 12× faster than SGS-SLAM.
@@
| GS4 (Ours) | 22.70 | 0.903 | 0.191 | 166k |
Zero-shot Experiments. In Table 6, we report quantitative zero-shot results. For NYUv2, the numbers are averaged over three scenes, and for TUM-RGBD, they are also averaged over three scenes. Per-scene results are provided in the appendix. On NYUv2, our method outperforms all other GS-based SLAM approaches across all rendering metrics, achieving a 15–20 relative improvement over the runner-up while using significantly fewer Gaussians. Qualitative comparisons are also provided in the appendix. On TUM-RGBD, our method outperforms the baselines in terms of SSIM and the number of Gaussians, and closely matches the best performance in other metrics, despite relying primarily on a feed-forward model trained on ScanNet.
Ablation Study. We conduct an ablation study using all ScanNet test scenes, as shown in Table 7. The results demonstrate that both the Gaussian Refinement Network and the Few-Iteration Joint Gaussian–Pose Optimization are critical to the performance of GS4. Additionally, Gaussian pruning significantly reduces the number of Gaussians without sacrificing accuracy.
| Design Choice | PSNR [dB] | SSIM | LPIPS | Depth L1 | mIoU | ATE | Gs Num |
| GS Prediction | 15.05 | 0.461 | 0.662 | 33.17 | 40.11 | 6.99 | 133k |
| + GS Refinement | 16.1 | 0.556 | 0.584 | 29.81 | 44.65 | 7.00 | 581k |
| + 1-Iter. Optimization | 22.72 | 0.851 | 0.337 | 6.24 | 63.58 | 6.97 | 666k |
| + GS Pruning (Full SLAM) | 22.62 | 0.851 | 0.338 | 6.46 | 63.71 | 6.97 | 321k |
Ablation Study of Few-Iteration Joint Gaussian–Pose Optimization. In Sec. 3.3.2, we propose to perform few-iteration joint Gaussian–pose optimization and feed the refined poses back into the tracking system to further improve pose accuracy. In Table 8, we ablate (i) Gaussian-only optimization and (ii) joint Gaussian–pose optimization with pose feedback into the tracking system. For the rendering metrics, joint Gaussian–pose optimization yields slight but consistent improvements over Gaussian-only optimization on both datasets. For tracking accuracy, we observe that pose feedback is particularly effective when high-quality ground truth is available, as in ScanNet++, whose RGB-D images are of much higher quality than those in ScanNet.
| Dataset | Optimization | PSNR | SSIM | LPIPS | GS Num | ATE RMSE |
| ScanNet | Gaussian | 24.39 | 0.881 | 0.302 | 225k | 6.94 |
| Gaussian–Pose | 24.55 | 0.885 | 0.299 | 224k | 6.95 | |
| ScanNet++ | Gaussian | 20.88 | 0.736 | 0.282 | 90k | 6.34 |
| Gaussian–Pose | 21.06 | 0.740 | 0.281 | 85k | 5.37 |
@@ 5 Conclusion
We present GS4, a novel SLAM system that incrementally constructs and updates a 3D semantic scene representation from a RGB-D video with a learned generalizable network. Our novel Gaussian refinement network and few-iteration joint Gaussian-pose optimization significantly improve the performance of our approach. Our experiments demonstrate state-of-the-art semantic SLAM performance on the ScanNet benchmark while running 10x faster and using 10x less Gaussians than baselines. The model also showed strong generalization capabilities through zero-shot transfer to the NYUv2 and TUM RGB-D datasets. In future work, we will further improve the computational speed of GS4 and explore options for a pure RGB-based SLAM approach.
@@ Appendix A More Experimental Setup
For ScanNet, we evaluate on six scenes—scene0000, scene0059, scene0106, scene0169, scene0181, and scene0207—which are commonly used by existing SLAM methods. For ScanNet++, we use six scenes in total: 8b5caf3398, 3f15a9266d, e7af285f7d, 99fa5c25e1, 09c1414f1b, and 9071e139d9. For NYUv2, we use the sequences bedroom0051out, diningroom0031out, and studentlounge0001out. For TUM RGB-D, we evaluate on rgbddatasetfreiburg1desk, rgbddatasetfreiburg2xyz, and rgbddatasetfreiburg3longofficehousehold.
Appendix B Ablation on semantic head
We present the semantic head ablation in Tab. 9. The versions with and without semantic prediction exhibit similar rendering performance, suggesting that semantics do not noticeably change reconstruction quality. This trend is consistent with the behavior of SGS-SLAM and SplaTAM. SGS-SLAM is built on top of SplaTAM by adding semantic colors to Gaussians and introducing a semantic loss for tracking and mapping. However, on the ScanNet, ScanNet++ and TUM RGB-D datasets, incorporating semantics actually degrades performance, with SGS-SLAM performing worse than SplaTAM. Taken together, these results indicate that incorporating semantic information does not inherently affect SLAM tracking or reconstruction quality. Nevertheless, we argue that having a unified backbone that jointly supports both SLAM and semantics remains valuable, as it allows a single model to deliver both accurate geometry and rich semantic understanding of the scene.
| Semantic Head | PSNR | SSIM | LPIPS | GS Num | ATE RMSE |
| False | 24.53 | 0.885 | 0.292 | 254k | 6.96 |
| True | 24.55 | 0.885 | 0.299 | 224k | 6.95 |
Appendix C More Results on GS4 with Different Prediction Stages
In the main paper, we use the second stage of the backbone as the default prediction stage. Here, we additionally report results using the third and fourth stages of the backbone as the prediction stage in Table 10 and Table 11. Using the fourth stage substantially reduces the number of Gaussians in the scene—by roughly fewer—while only slightly degrading the rendering quality and the semantic performance. As shown in Table 10, the fourth stage achieves PSNR that is comparable to the second stage, despite using only one tenth of the Gaussians, highlighting a favorable trade-off between reconstruction efficiency and fidelity. We also observe that, under the same prediction stage, using a lower input resolution leads to only a small drop in rendering performance, but quite significantly degraded semantic performance. This is because our model is ultimately trained at a higher resolution of , and downscaling the input reduces the amount of fine-grained semantic detail available per frame. Consequently, the per-frame semantic predictions become less accurate, which accumulates over the sequence and results in a noticeable drop in semantic performance at the scene level.
| Res | Method | PSNR | SSIM | LPIPS | GS Num | mIOU(2D) | ATE |
| Splat-SLAM | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | |
| GS4 (Ours, 2nd stg) | 24.55 | 0.885 | 0.299 | 224k | 62.10 | 6.95 | |
| GS4 (Ours, 3rd stg) | 24.56 | 0.875 | 0.344 | 69k | 60.34 | 6.94 | |
| GS4 (Ours, 4th stg) | 24.21 | 0.859 | 0.397 | 23k | 60.12 | 6.97 | |
| Splat-SLAM | 20.67 | 0.684 | 0.438 | 87k | ✗ | 7.60 | |
| GS4 (Ours, 2nd stg) | 22.43 | 0.921 | 0.196 | 124k | 52.85 | 6.94 | |
| GS4 (Ours, 3rd stg) | 24.24 | 0.911 | 0.245 | 37k | 47.55 | 6.93 | |
| GS4 (Ours, 4th stg) | 23.65 | 0.891 | 0.323 | 12k | 48.71 | 6.95 |
Appendix D Comparison with SLAM Methods Using Offline Post-Optimization Refinement
We omitted Splat-SLAM [sandstrom2024splat], MonoGS [Matsuki:Murai:etal:CVPR2024] and OVO-Gaussian-SLAM [martins2024ovo] from the main paper because these methods perform an additional 26k optimization iterations on the full map on the complete video after SLAM has finished. This heavy post-optimization step substantially boosts their final reconstruction quality (around dB PSNR), but is against the mentality of SLAM for generating online and incremental outputs. It is not directly comparable to our setting, where we report results of online SLAM without such long global refinement (roughly 5 minutes of extra optimization per scene). To ensure a fair comparison, we therefore disable these final refinement iterations on the full map and evaluate the baselines under a setting that matches ours more closely.
In Table 10, we report a comparison between our method and Splat-SLAM on ScanNet at two different resolutions, and . The higher resolution is included because all other baselines report rendering metrics at , and this is also the resolution used in our main experiments. However, Splat-SLAM fails under this setting on ScanNet, making a direct comparison at full resolution impossible. To still provide a fair and informative comparison, we additionally evaluate at the lower resolution of , which is the default configuration for Splat-SLAM on ScanNet. This allows us to compare performance at the resolution where Splat-SLAM is stable, while simultaneously highlighting that it does not robustly handle the higher-resolution setting used by other methods. At , our method consistently outperforms Splat-SLAM when using either the second or the fourth backbone stage for prediction. With the second stage, we surpass Splat-SLAM by 1.76 dB in PSNR (8.5), 0.237 in SSIM (34.6), and 0.242 in LPIPS (55.3). Even when using the fourth stage, we still outperform Splat-SLAM while using only 13.8 of its Gaussians, demonstrating a significantly better trade-off between quality and efficiency.
In Table 11, we report a comparison between our method and MonoGS, OVO-Gaussian-SLAM, and Splat-SLAM on the NYUv2 and TUM RGB-D datasets. MonoGS and OVO-Gaussian-SLAM share exactly the same reconstruction metrics because OVO-Gaussian-SLAM is built on top of the MonoGS system and simply adds semantic labels using CLIP features to predict a label for each Gaussian; both methods share the same mapping and tracking pipeline for RGB-D reconstruction. It is clear that, without the long global refinement stage, our method surpasses MonoGS, OVO-Gaussian-SLAM, and Splat-SLAM on most metrics on both NYUv2 and TUM RGB-D. In conclusion, our method outperforms these baselines that rely on post-optimization refinement on most metrics, while using substantially fewer Gaussians.
| Dataset | Method | PSNR | SSIM | LPIPS | GS Num | ATE |
| MonoGS | 12.88 | 0.505 | 0.550 | 110k | - - | |
| OVO-Gaussian-SLAM | 12.88 | 0.505 | 0.550 | 110k | - - | |
| Splat-SLAM | 23.30 | 0.754 | 0.332 | 85k | – | |
| NYUv2 | GS4 (Ours, 2nd stg) | 22.24 | 0.866 | 0.254 | 298k | - - |
| GS4 (Ours, 3rd stg) | 21.86 | 0.826 | 0.398 | 72k | – | |
| GS4 (Ours, 4th stg) | 20.94 | 0.766 | 0.530 | 18k | - - | |
| MonoGS | 17.78 | 0.718 | 0.315 | 37k | 1.52 | |
| OVO-Gaussian-SLAM | 17.78 | 0.718 | 0.315 | 37k | 1.52 | |
| Splat-SLAM | 22.84 | 0.780 | 0.287 | 62k | 1.1 | |
| TUM | GS4 (Ours, 2nd stg) | 22.70 | 0.903 | 0.191 | 166k | 1.41 |
| GS4 (Ours, 3rd stg) | 22.32 | 0.876 | 0.273 | 41k | 1.42 | |
| GS4 (Ours, 4th stg) | 21.23 | 0.833 | 0.392 | 12k | 1.42 |
Appendix E Do 3DGS-Based Methods Improve with a Stronger Tracking Module from GO-SLAM?
We adopt off-the-shelf tracking algorithms [zhang2023goslam, murai2025mast3r] in our system. Specifically, we use the GO-SLAM tracking module on ScanNet, NYUv2, and TUM RGB-D, and the Mast3r-SLAM tracking module on ScanNet++. To clarify whether other 3DGS-based methods could also benefit from improved tracking when equipped with a similar module, and to enable a more fair comparison with existing 3DGS-based SLAM systems, we compare against Splat-SLAM and MonoGS. Splat-SLAM adopts the MonoGS mapping pipeline but integrates it into an enhanced DROID-SLAM tracking framework, which is similar to our use of GO-SLAM-style tracking (which itself is derived from DROID-SLAM) on ScanNet, NYUv2, and TUM RGB-D.
We report the results of Splat-SLAM, MonoGS, and our method on these three datasets in Table 10 and Table 11. As expected, with a stronger tracking module, Splat-SLAM achieves noticeably better rendering performance than MonoGS, confirming that improved tracking alone can boost reconstruction quality for 3DGS-based systems. However, even under this stronger tracking setting, our method still substantially outperforms Splat-SLAM on most metrics, indicating that our backbone design and Gaussian prediction strategy provide additional gains beyond what can be achieved by improved tracking alone.
To further examine whether stronger tracking could also benefit semantic 3DGS-based SLAM, we attempted to inject GO-SLAM poses into GS3LAM by using GO-SLAM’s local pose estimates to initialize GS3LAM’s tracking. In practice, this integration turned out to be unstable: GS3LAM’s internal tracking and mapping pipeline is tightly coupled to its own pose-update dynamics, and the externally initialized poses conflicted with its optimization trajectory. This mismatch caused the estimated trajectory to drift abruptly, after which the system kept spawning large numbers of Gaussians to compensate for the growing misalignment. As a result, memory usage rapidly escalated and the run eventually terminated with an out-of-memory failure.
Appendix F Reconstruction Results on ScanNet